# Recovery of rational functions via Hankel pencil method and sensitivities of the poles

### v1

@@ 4 Recovery of rational functions via Hankel matrix pencil method and sensitivities of the poles
In this section, we describe the main ideas of our new method for the rational reconstruction and present analysis of the sensitivities of the poles.
@@ 4.1 Description of the method
$\widehat{r}_{k}:=\frac{1}{2\pi}\int\limits_{0}^{2\pi}r(\mathrm{e}^{\mathrm{i}\varphi})\mathrm{e}^{-\mathrm{i}k\varphi}\,\mathrm{d}\varphi=\frac{1}{2\pi\mathrm{i}}\int\limits_{\mathbb{D}}\frac{r(t)}{t^{k}}\,\frac{\mathrm{d}t}{t}.$
@@
Note that coefficients $\widehat{r}_{k}$ can be also computed applying the quadrature formulas on some grid of points. In that case, we will use the function values instead of Fourier coefficients as input data for the algorithm. This question was discussed in details in [19], where the author consider the random access model and the Matsubara model as a set of grid points for the evaluation of the transform $\widehat{r}_{k}$. In this paper, we are interested in the presentation of a modification of the method from [19] and in the explanation of the results of this method regarding recovery of noisy data with the use of sensitivities.
@@ 4.2 Sensitivities of poles
$\rho_{n}^{(-)}$ $=\frac{1}{\left|\gamma_{j_{n}}^{(-)}\right|}\left\|\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\,\boldsymbol{e}_{n}\right\|_{2}^{2}$
@@
$\boldsymbol{q}^{(n)}=\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n},\quad\boldsymbol{p}^{(n)}=\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n},$ (4.17)
@@
$(\boldsymbol{p}^{(n)})^{T}\,\mathbf{H}^{(+)}_{M_{2}}(0)\,\boldsymbol{q}^{(n)}$ $=\left(\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n}\right)^{T}\,\mathbf{H}^{(+)}_{M_{2}}(0)\,\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n}$
@@
$\rho_{n}^{(+)}=\frac{1}{|\gamma_{j_{n}}^{(+)}\,(z_{j_{n}}^{(+)})^{-2}|}\bigg{\|}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n}\bigg{\|}_{2}^{2}$
@@
$\|\boldsymbol{\rho}^{(-)}\|_{1}$ $\leq\zeta^{(-)}_{\max}\,\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{2}^{2}\,\|\mathbf{V}^{\dagger}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{\mathrm{F}}^{2}$ (4.22)
$\|\boldsymbol{\rho}^{(+)}\|_{1}$ $\leq\zeta^{(+)}_{\max}\,\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\|_{2}^{2}\,\|\mathbf{V}^{\dagger}_{M_{2}}(\boldsymbol{z}^{(+)})\|_{\mathrm{F}}^{2},$ (4.23)
where $\zeta_{\max}^{(-)}:=\max_{j=1,\ldots,M_{1}}\zeta_{j}^{(-)}$ and $\zeta_{\max}^{(+)}:=\max_{j=1,\ldots,M_{2}}\zeta_{j}^{(+)}$.
@@
$S_{n,m}^{(-)}=\left(S^{(n)}_{m}\right)^{(-)}:=$ $\frac{1}{\gamma_{j_{n}}^{(-)}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M_{1}\}}^{\min\{M_{1},m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\bigg{)}_{m-k,n}$
$-z_{j_{n}}^{(-)}\,\sum\limits_{k=\max\{1,m-M_{1}+1\}}^{\min\{M_{1},m\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\bigg{)}_{m-k+1,n}\bigg{)}$ (4.24)
@@
$S_{n,m}^{(+)}=\left(S^{(n)}_{m}\right)^{(+)}:=$ $-\frac{1}{\gamma_{j_{n}}^{(+)}\,z_{j_{n}}^{-2}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M_{2}\}}^{\min\{M_{2},m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\bigg{)}_{m-k,n}$
$-\frac{1}{z_{j_{n}}^{(+)}}\,\sum\limits_{k=\max\{1,m-M_{2}+1\}}^{\min\{M_{2},m\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\bigg{)}_{m-k+1,n}\bigg{)}.$ (4.25)
@@
5 Numerical experiments
In this section, we present a few numerical experiments. The algorithm is implemented in Matlab and use IEEE standard floating point arithmetic with double precision. We use the notations $\tilde{z}_{j}$ and $\tilde{\gamma}_{j}$ for the poles and the coefficients, respectively, reconstructed by the method considered in Subsection 4.1. We define the errors for the poles $z_{j}$ and the coefficients $\gamma_{j}$ by the following formulas
[…]
Note that in the case when we deal with noisy data, we assume that a number of poles $M$, as well as numbers of poles located inside and outside the unit circle $M_{1}$ and $M_{2}$ respectively are known.
@@ 5 Numerical experiments
We employ $2N$ (noisy) Fourier coefficients $\hat{r}_{-k}(1+n_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{j}^{(-)}$, and $2N$ noisy Fourier coefficients $\hat{r}_{k}(1+n_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{j}^{(+)}$. By $(n_{k})_{k=1}^{2N-1}$ we denote the random noise that is generated in Matlab by $\sigma$*randn(2N), i.e., we take multiplicative noise drawn from the standard normal distribution. We take $L=N$ and consider different values of $N$ and $\sigma$.
To compute sensitivities $\rho_{j}^{(-)}$, $\rho_{j}^{(+)}$ and ,$\eta_{j}^{(-)}$, $\eta_{j}^{(+)}$, we use formulas (4.13), (4.19) and (4.26), (4.27) respectively. Upper bounds for sensitivities $\rho_{j}^{(-)}$ and $\rho_{j}^{(+)}$ are computed via (4.14) and (4.20) respectively.
@@
Example 5.1.
Let a rational function $r(z)$ in (1.1) be given by parameters $M=2$ and
$\boldsymbol{z}=(-0.1,-2.1),\quad\boldsymbol{\gamma}=(0.5,0.5)$
[…]
For $\sigma=10^{-3}$ and $\sigma=10^{-1}$, we compute 10 iterations and present the results in Table 1. Since we are interested to investigate which pole is more sensitive to perturbations, we compute reconstruction errors for poles $z_{1}^{(-)}=-0.1$ and $z_{1}^{(+)}=-2.1$ separately. Results for recovery of the coefficient vector are presented in Table 2.
@@
Analyzing numerical results in Table 1, we conclude that for $\sigma=0.1$ our method still reconstructs location of poles quite good. On the other hand, using more (noisy) Fourier coefficients also gives us better recovery results. We notice that a pole $z_{1}^{(-)}=-0.1$ is less sensitive to perturbations than a pole $z_{1}^{(+)}=-2.1$. This numerical observation is also confirmed with sensitivities values. For a pole $z_{1}^{(-)}=-0.1$, we have $\rho_{1}^{(-)}=2\cdot 10^{-1}$ and $\eta_{1}^{(-)}=1.97$, and for a pole $z_{1}^{(+)}=-2.1$, we have $\rho_{1}^{(+)}=9.52\cdot 10^{-1}$ and $\eta_{1}^{(+)}=6.65$. Both unstructured and structured sensitivities are larger for a pole $z_{1}^{(+)}=-2.1$ located outside the unit circle. We compute also vectors of structured perturbations
$(\boldsymbol{\mathcal{S}}^{(1)}_{2})^{(-)}=(2.0\cdot 10^{-1},2.0)^{T}\text{ and }(\boldsymbol{\mathcal{S}}^{(1)}_{2})^{(+)}=(4.2,8.82)^{T}.$
That means that a pole $z_{1}^{(-)}=-0.1$ is the most sensitive with respect to perturbations of the Fourier coefficient $\hat{r}_{-2}$, and a pole $z_{1}^{(+)}=-2.1$ with respect to perturbation of the Fourier coefficient $\hat{r}_{2}$.
@@
Example 5.2.
respectively. We consider first recovery from the exact data, i.e for $\sigma=0$. For $N=8$, we get the following reconstruction errors
@@
In case of recovery from noisy data, we compute 10 iterations and present recovery result for each pole $z_{j}$ in Tables 3 and 4. Analyzing numerical results, we conclude that a pole $z_{1}^{(-)}=0.9$ is the most sensitive to perturbations. This conclusion is confirmed with sensitivities presented in Table 6. Results for recovery of the coefficient vector are presented in Table 5.
@@
Table 7: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.3 for $\sigma=10^{-7}$
$N=4$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
@@
Table 8: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.3 for $\sigma=10^{-5}$
$N=4$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
@@
Then we consider recovery from noisy data with $\sigma=10^{-7}$ and $\sigma=10^{-5}$. We compute 10 iterations and present the corresponding results in Tables 7 and 8. Analyzing numerical results, we can conclude that the pole $z_{2}^{(+)}=50$ is the most sensitive to perturbations. For $\sigma=10^{-5}$ the method does not reconstruct location of this pole. This numerical observation is also predicted by the sensitivities values presented in Table 9. Both unstructured $\rho_{2}^{(+)}=2.204\cdot 10^{3}$ and structured $\eta_{2}^{(+)}=1.577\cdot 10^{4}$ sensitivities are sufficiently larger than sensitivities of other poles. Note, that locations of poles $z_{1}^{(-)}=0.2$, $z_{2}^{(-)}=0.5$ and $z_{1}^{(+)}=2$ are reconstructed even for $\sigma=10^{-1}$. We present results regarding recovery of the coefficient vector in Table 10.
@@
$N=4$ $\sigma=10^{-7}$ $\sigma=10^{-5}$
min $7.38$e–$04$ $3.79$e–$02$
max $7.38$e–$03$ $6.95$e–$01$
average $4$e–$03$ $4.11$e–$01$
Table 11: $\ell_{2}$-norms and upper bounds of $\ell_{2}$-norms of vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ for Example 5.3
$N=4$ $2.86$e+$01$ $6.866$e+$01$ $2.204$e+$03$ $6.477$e+$03$
Table 12: $\ell_{1}$-norms and upper bounds of $\ell_{1}$-norms of vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ for Example 5.3
$N=4$ $4.01$e+$01$ $5.725$e+$01$ $2.209$e+$03$ $6.802$e+$03$
Further we compute the $\ell_{1}$- and $\ell_{2}$-norms of the vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ and their upper bounds obtained in (4.21), (4.22) and (4.23). We present the results in Tables 11 and 12. Analyzing these results, we conclude that for this example upper bounds (4.21), (4.22) and (4.23) are quite sharp.

### v2

@@ 4 Recovery of rational functions via Hankel matrix pencil method and sensitivities of the poles
In this section, we describe the main ideas of our new method for the recovery of a rational function of the form (1.1) and present the sensitivity analysis for its poles.
@@ 4.1 Description of the method
$\widehat{r}_{k}:=\frac{1}{2\pi}\int\limits_{0}^{2\pi}r(\mathrm{e}^{\mathrm{i}\varphi})\mathrm{e}^{-\mathrm{i}k\varphi}\,\mathrm{d}\varphi=\frac{1}{2\pi\mathrm{i}}\int\limits_{\mathbb{D}}\frac{r(t)}{t^{k}}\,\frac{\mathrm{d}t}{t}.$ (4.1)
@@
1. Note that coefficients $\hat{r}_{k}$ can be also computed applying the quadrature formulas on some grid of points. In that case, we will use the function values of $r$ instead of its Fourier coefficients as input data for the algorithm. Let us chose points $t_{n}=\mathrm{e}^{\frac{2\pi\mathrm{i}n}{4N}}$, $n=0,...,4N-1$ and let $(R_{k})_{k=0}^{4N-1}$ be the DFT of $(r(t_{n}))_{n=0}^{4N-1}$. Then $\hat{r}_{k}$ in (4.1) can be computed as $\hat{r}_{k}=R_{k+2N}$, $k=-2N,...,2N-1$. This question was discussed in details in [21], where the author consider not only the points on the unit circle, but also the Matsubara model as a set of grid points for the evaluation of the transform $\widehat{r}_{k}$. On the one hand, we are interested in the presentation of a modification of the method from [21] and in the explanation of the results of this method for the recovery from noisy data with the use of the sensitivities. In that case, we have to employ the given noisy coefficients $\widehat{r}_{k}$ as input data. On the other hand, we also want to compare the performance of known methods for rational approximation with our new approach based on Hankel pencil. To guarantee fair comparison, we assume to have given noisy values $(r(t_{n}))_{n=0}^{4N-1}$.
2. The method based on using the given coefficients $\hat{r}_{k}$, $k=-2N,...,2N$ has the overall computational cost of $\mathcal{O}(NL^{2})$, which is determined by computation of SVD of Hankel matrices $\mathbf{H}^{(-)}_{2N-L,L+1}$ and $\mathbf{H}^{(+)}_{2N-L,L+1}$. The complexity of the modified algorithm based on using the values $(r(t_{n}))_{n=0}^{4N-1}$ remains the same, since the additional operation computation of FFT of $(r(t_{n}))_{n=0}^{4N-1}$ requires only $\mathcal{O}(N\log N)$ flops.
@@ 4.2 Sensitivities of poles
$\rho_{n}^{(-)}$ $=\frac{1}{\left|\gamma_{j_{n}}^{(-)}\right|}\left\|\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\,\boldsymbol{e}_{n}\right\|_{2}^{2}$
@@
$\boldsymbol{q}^{(n)}=\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n},\quad\boldsymbol{p}^{(n)}=\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n},$ (4.18)
@@
$(\boldsymbol{p}^{(n)})^{T}\,\mathbf{H}^{(+)}_{M_{2}}(0)\,\boldsymbol{q}^{(n)}$ $=\left(\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n}\right)^{T}\,\mathbf{H}^{(+)}_{M_{2}}(0)\,\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n}$
@@
$\rho_{n}^{(+)}=\frac{1}{|\gamma_{j_{n}}^{(+)}\,(z_{j_{n}}^{(+)})^{-2}|}\bigg{\|}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n}\bigg{\|}_{2}^{2}$
@@
$\|\boldsymbol{\rho}^{(-)}\|_{1}$ $\leq\zeta^{(-)}_{\max}\,\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{2}^{2}\,\|\mathbf{V}^{-1}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{\mathrm{F}}^{2}$ (4.23)
$\|\boldsymbol{\rho}^{(+)}\|_{1}$ $\leq\zeta^{(+)}_{\max}\,\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\|_{2}^{2}\,\|\mathbf{V}^{-1}_{M_{2}}(\boldsymbol{z}^{(+)})\|_{\mathrm{F}}^{2},$ (4.24)
where $\zeta_{\max}^{(-)}:=\max_{j=1,\ldots,M_{1}}\zeta_{j}^{(-)}$ and $\zeta_{\max}^{(+)}:=\max_{j=1,\ldots,M_{2}}\zeta_{j}^{(+)}$. Naturally, arises the question how sharp are the obtained inequalities (4.224.24). In Example 5.3, we present the numerical computations for the exact values of the $\ell_{2}$- and $\ell_{1}$-norms of the vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ and their upper bound. Based on the results presented in Table 10, we can conclude that estimates (4.224.24) are quite sharp.
@@
$S_{n,m}^{(-)}=\left(S^{(n)}_{m}\right)^{(-)}:=$ $\frac{1}{\gamma_{j_{n}}^{(-)}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M_{1}\}}^{\min\{M_{1},m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg{)}_{m-k,n}$
$-z_{j_{n}}^{(-)}\,\sum\limits_{k=\max\{1,m-M_{1}+1\}}^{\min\{M_{1},m\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg{)}_{m-k+1,n}\bigg{)}$ (4.25)
@@
$S_{n,m}^{(+)}=\left(S^{(n)}_{m}\right)^{(+)}:=$ $-\frac{1}{\gamma_{j_{n}}^{(+)}\,z_{j_{n}}^{-2}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M_{2}\}}^{\min\{M_{2},m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg{)}_{m-k,n}$
$-\frac{1}{z_{j_{n}}^{(+)}}\,\sum\limits_{k=\max\{1,m-M_{2}+1\}}^{\min\{M_{2},m\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg{)}_{m-k+1,n}\bigg{)}.$ (4.26)
@@
5 Numerical experiments
In this section, we present a few numerical experiments. The algorithms are implemented in Matlab and use IEEE standard floating point arithmetic with double precision. By $\tilde{z}_{j}$ and $\tilde{\gamma}_{j}$ we denote reconstructed poles and coefficients, respectively, and by the following formulas
[…]
we define the corresponding reconstruction errors. Note that in the case when we deal with noisy data, we assume that a number of poles $M$, as well as numbers of poles located inside and outside the unit circle $M_{1}$ and $M_{2}$, respectively, are known.
In Examples 5.1-5.3, we demonstrate recovery results and analyze sensitivities of poles with the method developed in Section 4. We employ $2N$ (noisy) Fourier coefficients $\hat{r}_{-k}(1+\varepsilon_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{j}^{(-)}$, and $2N$ noisy Fourier coefficients $\hat{r}_{k}(1+\varepsilon_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{j}^{(+)}$. By $(\varepsilon_{k})_{k=1}^{2N}$ we denote the random noise that is generated in Matlab by $\sigma$*randn(2N), i.e., we take multiplicative noise drawn from the standard normal distribution. We take $L=N$ and consider different values of $N$ and $\sigma$. To compute sensitivities $\rho_{j}^{(-)}$, $\rho_{j}^{(+)}$ and ,$\eta_{j}^{(-)}$, $\eta_{j}^{(+)}$, we use formulas (4.14), (4.20) and (4.27), (4.28), respectively. Upper bounds for sensitivities $\rho_{j}^{(-)}$ and $\rho_{j}^{(+)}$ are computed via (4.15) and (4.21), respectively.
@@ 5 Numerical experiments
Let a rational function $r(z)$ in (1.1) be given by parameters $M=2$ and
$\boldsymbol{z}=(-0.1,-2.1),\quad\boldsymbol{\gamma}=(0.5,0.5)$
[…]
For $\sigma=10^{-3}$ and $\sigma=10^{-1}$, we compute 10 iterations and present the results in Table 1. Since we are interested to investigate which pole is more sensitive to perturbations, we compute reconstruction errors for poles $z_{1}^{(-)}=-0.1$ and $z_{1}^{(+)}=-2.1$ separately. Results for the recovery of the coefficient vector are presented in Table 2.
Table 7: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.3 for $N=4$
$\sigma=10^{-7}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
@@
$\sigma=10^{-5}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
@@
Analyzing numerical results in Table 1, we conclude that for $\sigma=0.1$ our method still reconstructs the locations of the poles quite good. On the other hand, using more (noisy) Fourier coefficients also gives us better recovery results. We notice that a pole $z_{1}^{(-)}=-0.1$ is less sensitive to perturbations than a pole $z_{1}^{(+)}=-2.1$. This numerical observation is also confirmed with sensitivities values. For a pole $z_{1}^{(-)}=-0.1$, we have $\rho_{1}^{(-)}=2\cdot 10^{-1}$ and $\eta_{1}^{(-)}=1.97$, and for a pole $z_{1}^{(+)}=-2.1$, we have $\rho_{1}^{(+)}=9.52\cdot 10^{-1}$ and $\eta_{1}^{(+)}=6.65$. Both unstructured and structured sensitivities are larger for a pole $z_{1}^{(+)}=-2.1$ located outside the unit circle. We compute also vectors of structured perturbations
$(\boldsymbol{\mathcal{S}}^{(1)}_{2})^{(-)}=(2.0\cdot 10^{-1},2.0)^{T}\text{ and }(\boldsymbol{\mathcal{S}}^{(1)}_{2})^{(+)}=(4.2,8.82)^{T},$
which means that a pole $z_{1}^{(-)}=-0.1$ is the most sensitive with respect to the perturbation of the Fourier coefficient $\hat{r}_{-2}$, and a pole $z_{1}^{(+)}=-2.1$ with respect to the perturbation of the Fourier coefficient $\hat{r}_{2}$.
@@
Example 5.2.
respectively. First, we consider recovery from the exact data, i.e for $\sigma=0$. For $N=8$, we get the following reconstruction errors
[…]
$N=4$ $\sigma=10^{-7}$ $\sigma=10^{-5}$
min $7.38$e–$04$ $3.79$e–$02$
max $7.38$e–$03$ $6.95$e–$01$
average $4$e–$03$ $4.11$e–$01$
In case of recovery from noisy data, we compute 10 iterations and present recovery result for each pole $z_{j}$ in Tables 3 and 4. Analyzing numerical results, we conclude that the pole $z_{1}^{(-)}=0.9$ is the most sensitive to perturbations. This conclusion is confirmed with the sensitivities presented in Table 6. Results for the recovery of the coefficient vector are presented in Table 5.
Table 10: $\ell_{p}$-norm and upper bound of $\ell_{p}$-norm of $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ for Example 5.3 for $p=1,2$
$N=4$ $2.86$e+$01$ $6.866$e+$01$ $2.204$e+$03$ $6.477$e+$03$
$N=4$ $4.01$e+$01$ $5.725$e+$01$ $2.209$e+$03$ $6.802$e+$03$
Example 5.3.
Then we consider recovery from noisy data with $\sigma=10^{-7}$ and $\sigma=10^{-5}$. We compute 10 iterations and present the corresponding results in Table 7. Analyzing numerical results, we can conclude that the pole $z_{2}^{(+)}=50$ is the most sensitive to perturbations. For $\sigma=10^{-5}$ the method does not reconstruct location of this pole correctly. This numerical observation is also predicted by the sensitivities values presented in Table 8. Both unstructured $\rho_{2}^{(+)}=2.204\cdot 10^{3}$ and structured $\eta_{2}^{(+)}=1.577\cdot 10^{4}$ sensitivities are sufficiently larger than sensitivities of other poles. Note, that the locations of the poles $z_{1}^{(-)}=0.2$, $z_{2}^{(-)}=0.5$ and $z_{1}^{(+)}=2$ are reconstructed even for $\sigma=10^{-1}$. We present results regarding recovery of the coefficient vector in Table 9.
@@
In Subsection 4.2, we derived the upper bounds (4.224.24) for the $\ell_{2}$- and the $\ell_{1}$-norms of vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ and we want to understand how sharp these estimates are. In order to do it, we compute the exact values for the corresponding norms and the upper bounds obtained in (4.224.24) and present the corresponding results in Table 10. Comparing these results we can conclude that the estimates (4.224.24) are quite sharp.
Table 11: Reconstruction results for $r$ from Example 5.1
Method via Hankel pencil AAA method
@@
min $1.52$e–$08$ $1.23$e–$08$ min $3.54$e–$08$ $4.55$e–$08$
max $1.19$e–$07$ $5.51$e–$08$ max $6.62$e–$07$ $4.84$e–$07$
average $7.51$e–$08$ $2.89$e–$08$ average $3.83$e–$07$ $2.82$e–$07$
Method via Hankel pencil AAA method
[…]
min $4.34$e–$03$ $1.06$e–$03$ min $2.37$e–$02$ $2.27$e–$02$
max $1.23$e–$02$ $4.24$e–$03$ max $2.05$e+$00$ $7.56$e–$02$
average $7.64$e–$03$ $2.71$e–$03$ average $4.52$e–$01$ $4.55$e–$02$
Table 12: Reconstruction results for $r$ from Example 5.2
Method via Hankel pencil AAA method
[…]
min $1.96$e–$05$ $4.79$e–$06$ min $7.03$e–$05$ $7.18$e–$05$
max $1.37$e–$04$ $6.22$e–$05$ max $4.49$e–$04$ $3.43$e–$04$
average $5.89$e–$05$ $2.34$e–$05$ average $2.98$e–$04$ $2.25$e–$04$
Method via Hankel pencil AAA method
[…]
min $2.2$e–$03$ $2.42$e–$03$ min $1.54$e–$02$ $9.73$e–$03$
max $1.37$e–$02$ $4.86$e–$02$ max $6.69$e–$02$ $3.54$e–$02$
average $6.93$e–$03$ $1.06$e–$02$ average $4.42$e–$02$ $5.75$e–$02$
In Example 5.4 we present the comparison results of the performance of our approach with the AAA method for rational approximation. We employ $4N$ (noisy) function values $r(t_{k})(1+\varepsilon_{k})$ with $t_{k}=\mathrm{e}^{\frac{2\pi\mathrm{i}k}{4N}}$ for $k=0,\ldots,4N-1$, as input information for both methods. By $(\varepsilon_{k})_{k=0}^{4N-1}$ as above we denote the random noise that is generated in Matlab by $\sigma$*randn(4N), i.e., we take multiplicative noise drawn from the standard normal distribution. The corresponding modification of our method is discussed in Remark 4.1. We referee the reader to [14] regarding description of the AAA method. Note only that AAA takes the form of iteration with respect to a type of a rational approximant that is represented at each step in the barycentric form. In case of reconstruction of rational function of type $(M-1,M)$ using exact data, AAA terminates after $M+1$ iteration steps and returns $r$ as
$r(t)=\sum\limits_{s=1}^{M+1}\frac{r(t_{k_{s}})w_{s}}{t-t_{k_{s}}}\bigg{/}\sum\limits_{s=1}^{M+1}\frac{w_{s}}{t-t_{k_{s}}},$
where points $t_{k_{s}}$ are chosen by some greedy algorithm and weights $w_{s}$ are computed by solving a linear least-squares problem over the subset of sample points that have not been selected as support points, $\{t_{0},\ldots,t_{4N-1}\}\setminus\{t_{k_{1}},\ldots,t_{k_{M+1}}\}$ (see also [8, 9, 10] for details). The AAA method has the overall computational cost of $\mathcal{O}(NM^{3})$.
Table 13: Reconstruction results for $r$ from Example 5.3
Method via Hankel pencil AAA method
[…]
min $4.35$e–$03$ $7.99$e–$05$ min $1.26$e–$01$ $4.89$e–$03$
max $3.00$e–$02$ $5.66$e–$04$ max $2.03$e+$00$ $7.86$e–$02$
average $1.42$e–$02$ $2.66$e–$04$ average $1.48$e+$00$ $4.44$e–$02$
Method via Hankel pencil AAA method
[…]
min $7.89$e–$04$ $4.48$e–$06$ min $2.6$e–$02$ $1.01$e–$03$
max $1.28$e–$03$ $5.58$e–$05$ max $2.6$e–$01$ $1.01$e–$02$
average $2.98$e–$03$ $2.4$e–$05$ average $1.25$e–$01$ $4.84$e–$03$
Example 5.4.
We consider again recovery of rational functions $r$ from Examples 5.1-5.3. We reconstruct $r$ from $4N$ samples $r(t_{k})(1+\varepsilon_{k})$, $k=0,\ldots,4N-1$ by using our method based on Hankel pencil and AAA. In Tables 11- 13 we present the results for $N=15$, $L=20$ and different values of $\sigma$. We see that for the rational function from Example 5.2 both methods work very well.
Analysing results of Tables 11 and 13, we conclude that for $r$ from Examples 5.1 and 5.3 we get a slightly better recovery from noisy input data with our method. The reason for this is that the method based on the Hankel pencil works better for the recovery of poles with certain locations. For Example 5.3 the poles $z_{1}^{(-)}=0.2$, $z_{2}^{(-)}=0.5$ and $z_{1}^{(+)}=2$ are reconstructed equally good by both methods, while the Hankel pencil based approach shows better performance for the recovery of the pole $z_{2}^{(+)}=50$ than the AAA method (see Tables 14).
Table 14: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.3 with Hankel pencil and the AAA methods
Method via Hankel pencil
$\sigma=10^{-7}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
min $2.82$e–$08$ $6.98$e–$08$ $2.08$e–$07$ $4.35$e–$03$
max $5.09$e–$07$ $3.92$e–$07$ $1.14$e–$06$ $3$e–$02$
average $2.51$e–$07$ $1.79$e–$07$ $2.08$e–$07$ $1.72$e–$02$
AAA method
$\sigma=10^{-7}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
min $1.15$e–$07$ $6.42$e–$08$ $6.65$e–$07$ $1.26$e–$01$
max $7.4$e–$07$ $5.24$e–$07$ $2.55$e–$06$ $2.03$e+$00$
average $4.63$e–$07$ $2.34$e–$07$ $1.19$e–$06$ $1.14$e+$00$